# Using Anomaly Detection to Detect Poisoning Attacks in Federated Learning Applications

### v3

@@ 4.2 Performance Evaluation
Table III shows the poisoning rate identified by the proposed framework for each edge in the federated setting for HAR. Here Edge 1 and Edge 2 are benign devices and Edge 3 is the malicious device that launches different data poisoning attacks in each global round, whereas the $P$ value in the table shows the threshold, which is calculated using $h_{\text{test}}$ and $\sigma$. Any updates ($W_{i}$) from an edge device for which it’s corresponding $DA_{i}$ has a value of $h_{i}$ greater than $P$ will be marked as malicious or poisoned and removed from global aggregation. It can be seen that in Table III, $h_{1}$ and $h_{2}$ for Edge 1 and Edge 2, respectively, have a value smaller than $P$. Hence, the updates $W_{1}$ and $W_{2}$ will be included in global aggregation, while $W_{3}$ will be excluded from global aggregation as its corresponding $h_{3}$ is greater than $P$. It can be seen that our proposed framework can not only detect the attacks but also provide insights into the percentage (a general idea about possible poison percentage) of poisoned data being used to train a malicious local model. For example, for LS we swapped the label of some classes (for ECG classification two classes and for HAR 10 classes) while keeping the rest of the labels in their original form, thereafter it gives a value of 47.7 for ECG and 81.2 for HAR. Similarly, Tables IV, VI and V present the performance of our proposed framework to detect data poisoning attacks on ECG classification, model poisoning attacks on ECG classification, and model poisoning attacks on HAR, respectively. Hence, these results show that the proposed framework can identify malicious updates which are then removed from the global aggregation.
TABLE III: Detection of data poisoning attacks on HAR
@@
TABLE IV: Detection of data poisoning attacks on ECG classification
@@
TABLE V: Detection of model poisoning attacks on HAR classification
@@
TABLE VI: Detection of model poisoning attacks on ECG classification
@@ 5 Comparison
Table VII shows a comparison of our proposed framework with some of the SOTA methods for detecting poisioning attacks in FL [21, 33, 27, 4, 34, 24, 25, 35, 36, 28, 37, 22], where the Category column presents the type of detection mechanism, attack type presents type of attack, the Model Accuracy column presents the accuracy of global model under attack compared with the accuracy of global model without any attack, the Data Distribution column presents the type of data distribution among clients: independent identically distributed (IID), non identically distributed (Non-IID), $d$ represents model size (depth) of each client and $K$ is the total number of edge devices/clients. It can be seen that our proposed framework provide more desirable features than all other SOTA methods. Our proposed framework does not need knowledge about the number of attackers. It can detect both types of attacks, i.e., model and data poisoning attacks. Moreover, its time complexity is the least compared to others. Additionally, the proposed method can be used in both IID and non-IID data distribution with high accuracy.
TABLE VII: Comparison with SOTA methods.
@@ 6 Further Discussions
In this section, we provide some further discussions including limitations of our proposed framework and future work.
First of all, we would like to highlight that the core of our proposed method is an existing public dataset for the target machine learning model our FL system tries to collective build. The public dataset is considered validated and cannot be compromised by the attacker. While this assumption can be considered quite strong, it is reasonable for many real-world applications given the richness of public and open datasets in many fields. The public dataset effectively provides a “ground truth” benchmark so that we can train a reference model (RM) and an auditor model to test gradients (and updates) sent by all clients of the FL system. Note that such a public dataset does not need to stay static, but can keep evolving by continuously accepting new and validated samples, like how many open datasets are being maintained in many fields. Similarly, mistakenly added malicious samples can also be removed from the public dataset. In a way, our proposed method leverages the community-wide review processes to maintain the integrity of the public dataset.
Although our proposed method has a number of merits, it also has some limitations. We discuss some of such limitations below, together with corresponding future research directions.
@@
    The need to maintain a stable and validated public dataset: The dependence on a well-validated public dataset means that our method will not work if drastically different new data samples and even new classes keep emerging. Examples include automatic detection of rare diseases or rapidly changing viruses (e.g., COVID or flu viruses), and intrusion detection systems where attackers keep changing their attack behaviors. In other words, our proposed method is suitable only for applications where 1) class labels are stable and representative samples of each class do not change too rapidly, or 2) a public dataset can be easily maintained by an active community despite the fast evolution of data samples. In the future, we will study how to relax the dependency on the public dataset, which likely requires new ideas to update the RM, e.g., based on some distributed consensus protocols.
@@
    Dependence on a trusted third party or component: Conceptually speaking, our proposed model relies on a trusted third party or a trusted component of the global server – the auditor model + the reference model + the public dataset. Such dependencies are widely used in cryptography. In some applications, such a trusted party or model may be difficult to found or established, e.g., when the clients do not have trust on each other and cannot agree on anything they can trust collectively. If it is possible to remove the trusted party/component or constructed a similar entity under a “zero trust” environment is an interesting future research direction.
@@
    Sensitivity to “small” poison in updates: The anomaly detection component of our proposed method still relies on a threshold to detect malicious clients, therefore it will not be able to detect “small” poisoning attacks. While in this case a single attacker may find it difficult to poison the global model, carefully coordinated colluding attacks of many malicious clients may be able to collectively make a difference. Detecting such small colluding poisoning attacks requires different approaches, and will be one interesting future research directions.
@@
    Limited experimental results: Although we conducted many experiments, the existence of many poisoning attacks and different configurations of the FL system means that our experimental results do not cover all aspects of our proposed method. For instance, we used only three edge devices and one attacker, and therefore did not consider different combinations of multiple attackers who could launch different attacks with different parameters and some of them could collude. In the future, we plan to conduct a more comprehensive set of experiments to investigate more aspects of our proposed method’s performance.
@@ 7 Conclusions
In this paper, we present a novel framework to detect poisoning attacks in FL applications. Our proposed framework can efficiently detect SOTA data and model poisoning attacks by observing the activations of the shared weights of the local models. Unlike most of the existing methods, our proposed framework can detect poisoning attacks without degrading the global model’s performance. In addition, while most of the existing methods need knowledge about the number of attackers in the network, which can limit their applications, our proposed method can detect poisoning attacks without any knowledge about the number of attackers in the network. Moreover, in principle, our proposed method can detect any number of attackers, especially if they do not collude, thanks to the use of the auditor model based on the public dataset, while many SOTA methods can only work up to a particular number of attackers. Additionally, the time complexity of our proposed framework is dependent only on the number of edge devices, with makes it suitable to be used with any size (in terms of depth) of the network, whereas the time complexity of existing SOTA methods depends on the number of edge devices as well as the size of the network. We tested our proposed framework under four different data poisoning attacks and four different model poisoning attacks for two healthcare applications, showing that the proposed framework could efficiently detect malicious updates and can exclude them from the global aggregation.

### v4

@@ 4.2 Performance Evaluation
Table 3 shows the poisoning rate identified by the proposed framework for each edge in the federated setting for HAR. Here Edge 1 and Edge 2 are benign devices and Edge 3 is the malicious device that launches different data poisoning attacks in each global round, whereas the $P$ value in the table shows the threshold, which is calculated using $h_{\text{test}}$ and $\sigma$. Any updates ($W_{i}$) from an edge device for which its corresponding $DA_{i}$ has a value of $h_{i}$ greater than $P$ will be marked as malicious or poisoned and removed from global aggregation. It can be seen that in Table 3, $h_{1}$ and $h_{2}$ for Edge 1 and Edge 2, respectively, have a value smaller than $P$. Hence, the updates $W_{1}$ and $W_{2}$ will be included in global aggregation, while $W_{3}$ will be excluded from global aggregation as its corresponding $h_{3}$ is greater than $P$. It can be seen that our proposed framework can not only detect the attacks but also provide insights into the percentage (a general idea about possible poison percentage) of poisoned data being used to train a malicious local model. For example, for LS, we swapped the labels of some classes (for ECG classification two classes and HAR 10 classes) while keeping the rest of the labels in their original form; thereafter, it gives a value of 47.7 for ECG and 81.2 for HAR. Similarly, Tables 4, 6 and 5 present the performance of our proposed framework to detect data poisoning attacks on ECG classification, model poisoning attacks on ECG classification, and model poisoning attacks on HAR, respectively. Hence, these results show that the proposed framework can identify malicious updates which are then removed from the global aggregation.
Table 3: Detection of data poisoning attacks on HAR
@@
Table 4: Detection of data poisoning attacks on ECG classification
@@
Table 5: Detection of model poisoning attacks on HAR classification
@@
Table 6: Detection of model poisoning attacks on ECG classification
@@ 5 Comparison
Table 7 shows a comparison of our proposed framework with some of the SOTA methods for detecting poisoning attacks in FL [21, 33, 27, 4, 34, 24, 25, 35, 36, 28, 37, 22], where the Category column presents the type of detection mechanism, attack type presents a type of attack, the Model Accuracy column presents the accuracy of global model under attack compared with the accuracy of the global model without any attack, the Data Distribution column presents the type of data distribution among clients: independent identically distributed (IID), non identically distributed (Non-IID), $d$ represents model size (depth) of each client and $K$ is the total number of edge devices/clients. It can be seen that our proposed framework provides more desirable features than all other SOTA methods. Our proposed framework does not need knowledge about the number of attackers. It can detect both types of attacks, i.e., model and data poisoning attacks. Moreover, its time complexity is the least compared to others. Additionally, the proposed method can be used in both IID and non-IID data distribution with high accuracy.
Table 7: Comparison with SOTA methods.
@@ 6 Further Discussions
In this section, we provide some further discussion, including the limitations of our proposed framework and future work.
First of all, we would like to highlight that the core of our proposed method is an existing public dataset for the target machine learning model our FL system tries to build collectively. The public dataset is considered validated and cannot be compromised by the attacker. While this assumption can be considered quite strong, it is reasonable for many real-world applications given the richness of public and open datasets in many fields. The public dataset effectively provides a “ground truth” benchmark so that we can train a reference model (RM) and an auditor model to test gradients (and updates) sent by all clients of the FL system. Note that such a public dataset does not need to stay static but can keep evolving by continuously accepting new and validated samples, such as how many open datasets are being maintained in many fields. Similarly, mistakenly added malicious samples can be removed from the public dataset. In a way, our proposed method leverages the community-wide review processes to maintain the integrity of the public dataset.
Although our proposed method has many merits, it also has some limitations. We discuss some of such limitations below, together with corresponding future research directions.
@@
    The need to maintain a stable and validated public dataset: The dependence on a well-validated public dataset means that our method will not work if drastically different new data samples and even new classes keep emerging. Examples include automatic detection of rare diseases or rapidly changing viruses (e.g., COVID or flu viruses), and intrusion detection systems where attackers keep changing their attack behaviors. In other words, our proposed method is suitable only for applications where 1) class labels are stable and representative samples of each class do not change too rapidly, or 2) an active community can easily maintain a public dataset despite the fast evolution of data samples. In the future, we will study how to relax the dependency on the public dataset, which likely requires new ideas to update the RM, e.g., based on some distributed consensus protocols.
@@
    Dependence on a trusted third party or component: Conceptually speaking, our proposed model relies on a trusted third party or a trusted component of the global server – the auditor model + the reference model + the public dataset. Such dependencies are widely used in cryptography. In some applications, such a trusted party or model may be difficult to find or establish, e.g., when the clients do not have trust in each other and cannot agree on anything they can trust collectively. Whether it is possible to remove the trusted party/component or construct a similar entity under a “zero trust” environment is an interesting future research direction.
@@
    Sensitivity to “small” poison in updates: The anomaly detection component of our proposed method still relies on a threshold to detect malicious clients, therefore it will not be able to detect “small” poisoning attacks. While in this case, a single attacker may find it difficult to poison the global model, carefully coordinated colluding attacks of many malicious clients may be able to collectively make a difference. Detecting such small colluding poisoning attacks requires different approaches and will be one interesting future research direction.
@@
    Limited experimental results: Although we conducted many experiments, the existence of many poisoning attacks and different configurations of the FL system means that our experimental results do not cover all aspects of our proposed method. For instance, we used only three edge devices and one attacker and therefore did not consider different combinations of multiple attackers who could launch different attacks with different parameters, and some of them could collude. In the future, we plan to conduct a more comprehensive set of experiments to investigate more aspects of our proposed method’s performance.
@@ 7 Conclusions
In this paper, we present a novel framework to detect poisoning attacks in FL applications. Our proposed framework can efficiently detect SOTA data and model poisoning attacks by observing the activations of the shared weights of the local models. Unlike most of the existing methods, our proposed framework can detect poisoning attacks without degrading the global model’s performance. In addition, while most of the existing methods need knowledge about the number of attackers in the network, which can limit their applications, our proposed method can detect poisoning attacks without any knowledge about the number of attackers in the network. Moreover, in principle, our proposed method can detect any number of attackers, especially if they do not collude, thanks to the use of the auditor model based on the public dataset, while many SOTA methods can only work up to a particular number of attackers. Additionally, the time complexity of our proposed framework is dependent only on the number of edge devices, which makes it suitable to be used with any size (in terms of depth) of the network, whereas the time complexity of existing SOTA methods depends on the number of edge devices as well as the size of the network. We tested our proposed framework under four different data poisoning attacks and four different model poisoning attacks for two healthcare applications, showing that the proposed framework could efficiently detect malicious updates and exclude them from the global aggregation.